# www.entalus.com

### v1

@@
Abstract
We study solutions to systems of stream inclusions $f\in T(f)$, where $T$ is assumed to be causal in the sense that elements in output streams are determined by a finite history of inputs. For solving these inclusions we develop a correspondence of causality and contraction with respect to the prefix distance on streams. Now, based on this causality-contraction correspondence, we apply fixpoint principles for the spherically complete ultrametric space of streams to obtain solutions for causal stream inclusions. The underlying fixpoint iterations induce fixpoint induction principles for reasoning about solutions of causal stream inclusions. In addition, these fixpoint approximations induce anytime algorithms for computing finite stream prefixes of solutions. We illustrate the use of these developments for some central concepts of system design.
Keywords:
Formal Methods Fixpoint Approximation Systems Engineering.
Preprint
$23^{rd}$ Oct 2023
1 Introduction
@@
where $f\,\coloneqq\,(f_{1},\ldots,f_{n})$, for $n\in\mathbb{N}$, is a vector of infinite streams $f_{i}$ of values, and the vector- and multivalued stream transformer $T$ is causal in that every element in the output stream is completely determined by a finite history of inputs. These stream inclusions are ubiquitous in computer science and other fields of knowledge such as biology, economics, and artificial intelligence for modeling the evolution of systems under uncertainty. Typical sources of uncertainty hereby are lack of epistemic knowledge about the behavior of the system under consideration, as well as modeling artefacts such as nondeterminism and underspecification.
@@
If the stream transformer $T$ is deterministic then the stream inclusion (1) reduces to a system of stream equalities $f=T(f)$. Interacting systems of deterministic stream processors, for instance, are modeled as the least fixpoint of this equality, where $T$ is a Scott-continuous stream transformer in a complete partial order [21]. But this traditional denotational approach does not extend naïvely to unbounded nondeterminism [3, 60, 47, 45, 59, 8, 10].
@@
We are pursuing a different route in that we are investigating conditions under which fixpoints for causal stream transformers with unbounded nondeterminism, which is naturally modeled by means of multivalued maps, exist. We hereby distinguish between weakly and strongly causal stream transformers, whereby the latter notion additionally implies a strict, bounded delay for which outputs are also determined. Our main results for solving stream inclusions (1) are as follows.
@@
  • (Theorem 7.1) For strongly causal stream transformers $T$ with nonempty, compact codomains, solutions $f\in T(f)$ are contained in the (unique) fixpoint $F={\mathrm{sp}}_{T}(F)$ for the strongest post of $T$. This latter fixpoint is obtained as the limit of a Picard iteration with a bound on the approximation of each iteration.
@@
  • (Theorem 7.2) $f\in T(f)$ has a solution if $T$ is strongly causal, not identically empty, and all the codomains of $T$ are closed (in the topology induced by the prefix metric on streams). In addition, we identify a slightly stronger condition than strong causality for establishing the uniqueness of solutions of the stream inclusion (1). A fixpoint induction principle is provided for reasoning about these solutions.
  • (Theorem 7.3) $f\in T(f)$ has a solution if $T$ is strongly causal and all codomains of $T$ are nonempty and compact (in the topology induced by the prefix metric on streams).
@@
  • (Theorem 7.4) If the stream transformer $T$ is weakly causal then either $f\in T(f)$ has a solution or there exists a ball of streams with positive radius on which the distance, as measured by the prefix metric on streams, between input stream and corresponding output sets with respect to $T$ is, in a sense to be made more precise, invariant.
@@
Our developments are heavily based on the correspondence of logic-based notions of causality with metric-based notions of contraction (and non-expansion). More precisely, we show that a vector- and multivalued stream transformer is weakly causal if and only if it is non-expansive, and it is strongly causal if and only if it is contractive for the prefix metric. This approach is motivated by recent work of Broy [10] on stream-based system design calculus for strongly causal functionals, which is based on the existence of unique fixpoints for strongly causal, and therefore contracting functionals for the prefix metric on streams, by applying Banach’s fixpoint principle. Altogether, we proceed by:
  1. 1.
    Modeling nondeterministic and mutually dependent system components as conjunctions of causal stream inclusions;
  2. 2.
    Establishing the equivalence of vector- and multivalued causal stream transformers with contraction in a spherically complete ultrametric space of streams based on the prefix distance of streams;
  3. 3.
    Applying multivalued fixpoint principles in this ultrametric stream space for obtaining solutions of causal stream inclusions.
@@
As a starting point, in Section 3, the prefix distance between two streams is measured in terms of the longest common prefix, thereby obtaining a spherically complete ultrametric space of streams. This notion of distance, which derives from Cantor sets, generalizes to arbitrary stream products and also to sets of streams via the induced Hausdorff distance. Next, Section 4 reviews the familiar notions of stream transformers together with some composition operators, refinement, and contractual specification as the basis of a more general framework for systematically constructing stream transformers and their realization on computational machinery. The focus in Section 5 is on causal stream transformers, whose outputs are determined by a finite history of inputs. For instance, causality is preserved under composition and refinement of stream transformers.
@@
In Section 6 we develop a metric-based characterization of causal stream transformers. More specifically, we show that a nondeterministic stream transformer is causal if and only if it is contractive with respect to the given ultrametric on streams. This correspondence extends to a notion of Lipschitz contraction based on the Hausdorff distance between sets of streams, since stream transformers with non-empty compact images are Lipschitz contractive if and only if they are causal.
@@
Since the induced Hausdorff metric is a complete metric space on the set of non-empty compact sets, we obtain in Section 7 unique fixpoints for causal set transformers such as the weakest pre and the strongest post set transformers. Moreover, there is a linearly, strictly decreasing upper bound of the prefix distance between this fixpoint and its approximation by the underlying Picard iteration, which suggests an anytime approximation algorithm, we formulate induction principles in the spirit of Park’s lemma for fixpoints of set transformers, and we show that the set of fixpoints of a multivalued stream transformer $T$ is contained, and possibly strictly so, in the fixpoint of the strongest post transformer for $T$.
@@
Additional fixpoint results in Section 7 construct solutions for the stream inclusion (1) by showing that every strongly causal vector- and multivalued stream transformer has a fixpoint, as long as it is not identically empty and its codomain is restricted to closed sets only. We also identify a slightly stronger condition for which this fixpoint is unique. The underlying Picard iteration enables us to derive an induction principle for reasoning about these fixpoints. In Section 7 we also state some immediate consequences of the correspondence of causality and Lipschitz contractivity for fixpoints of weakly causal maps. Section 8 discusses futher consequences of the causality-contraction correspondence extensions, and Section 9 concludes with an outlook on the relevance of these developments for system design.
2 Preliminaries
@@
We are summarizing a smorgasbord of concepts and notation for topological and metric spaces, as used in the remainder of this treatise. For a metric space $(M,d)$ the sets $\mathrm{B}({x},{r})\,\coloneqq\,\{{y}\,\mid\,{d(x,y)<r}\}$ and $\mathrm{B}[{x},{r}]\,\coloneqq\,\{{y}\,\mid\,{d(x,y)\leq r}\}$ are called the open and closed balls of center $x$ and radius $r$, respectively. The family of open balls forms a base of neighborhoods for a uniquely determined Hausdorff topology on $M$, which is the topology induced by $d$ (on $M$). Open, closed, bounded, (dis)connected, convex, totally bounded (precompact), and compact sets are defined with respect to the metric-induced topology. The set of nonempty, closed, and bounded subsets of $M$, in particular, is denoted by ${\mathrm{CB}({M})}$, and ${\mathrm{Comp}({M})}$ is the set of nonempty compact subsets of $M$.
@@
The distance $d(a,B)$ of an element $a\in M$ to a non-empty set $B\subseteq M$ is defined by $d(a,B)=\inf_{b\in B}{d(a,b)}$. Clearly, $d(a,b)=d(a,\{b\})$ for all $a,b\in M$. For a bounded metric space $(M,d)$ the Hausdorff distance $\mathcal{H}_{d}(A,B)$ between two nonempty sets $A,B\subseteq M$ measures the ”longest path” to get from $A$ to $B$, or vice versa, from $B$ to $A$.
$\mathcal{H}_{d}(A,B)$ $\,\coloneqq\,\max{(\sup_{x\in A}{(d(x,B))},\sup_{y\in B}{(d(y,A))})}\mbox{\@.}$ (2)
@@
A sequence $(x_{k})_{k\in\mathbb{N}}$ in $M$ is Cauchy if and only if for all $\varepsilon>0$ there exists $N\in\mathbb{N}$ such that $d(x_{n},x_{m})<\varepsilon$ for all $n,m\geq N$. Such a sequence $x_{k}$ converges to $x^{*}\in M$ if and only if for each neighborhood $U$ of $x$ there exists $N\in\mathbb{N}$ such that $x_{n}\in U$ for all $n\geq N$. Now, the space $M$ is Cauchy complete if every Cauchy sequence converges to some $x^{*}\in M$. Equivalently, $M$ is Cauchy complete if and only if the intersection of nested sequences of closed balls whose radius approaches to 00 are non-empty.
@@
A map $T:M\to M$ (in a metric space $M$) is contractive if there exists a constant $l$ with $10<l<$ 1 such that $d(T(x),T(y))\leq l\cdot d(x,y)$ for all $x,y\in M$. Traditionally, Banach’s contraction principle establishes that in a Cauchy complete metric space $(M,d)$ there is a unique fixpoint for every contracting $T:M\to M$ [29]. Starting with the Picard iteration $x_{n+1}=T(x_{n})$ for $n\in\mathbb{N}$ with $x_{0}$ arbitrary, one concludes from the contraction property with Lipschitz constant $10\leq l<$ 1, that $d(x_{n+1},x_{n})\leq l\cdot d(x_{n},x_{n-1})$, and, therefore, $d(x_{n+1},x_{n})\leq\nicefrac{{l^{n}}}{{1-l}}\cdot d(x_{0},x_{1})$. From this, one concludes that the sequence $(x_{n})_{n\in\mathbb{N}}$ is Cauchy, and, for completeness, that its limit $x^{*}\in M$ is a fixpoint. This fixpoint is unique, since we get $x^{*}=y^{*}$ for any two fixpoints from $d(x^{*},y^{*})=d(T(x^{*}),T(y^{*}))\leq l\cdot d(x^{*},y^{*})$.
@@
A map $T:M\to M$ is said to be shrinking if $d(T(x),T(y))<d(x,y)$ for all $x,y\in M$. A shrinking map $T$ need not have a fixed point in a complete metric space.
Example 1

### v2

# Solving Causal Stream Inclusions

@@
Abstract.
We study solutions to systems of stream inclusions of the form $f\in T(f)$, where the nondeterministic transformer $T$ on $\omega$-infinite streams is assumed to be causal in the sense that elements in output streams are determined by a finite prefix of inputs. We first establish a correspondence between logic-based causality and metric-based contraction. Based on this causality-contraction connection we then apply fixpoint principles to the spherically complete ultrametric space of streams to construct solutions of stream inclusions. The underlying fixpoint iterations induce fixpoint induction principles to reason about these solutions. In addition, the fixpoint approximation provides an anytime algorithm with which finite prefixes of solutions can be calculated. These developments are illustrated for some central concepts of system design.
1. Introduction
@@
where $f\,\coloneqq\,(f_{1},\ldots,f_{n})$, for $n\in\mathbb{N}$, is a vector of infinite streams $f_{i}$ of values, and the vector- and multivalued stream transformer $T$ is causal in that every element in the output stream is completely determined by a finite history of inputs. These stream inclusions are ubiquitous in computer science and other fields of knowledge such as biology, economics, and artificial intelligence to model the evolution of dynamic systems under uncertainty.
Typical sources of uncertainty are lack of epistemic knowledge about the behavior of the system under consideration as well as modeling artifacts such as underspecification, whereby a system is said to be nondeterministic and a system specification is underspecified if for some input several outputs are admitted. Underspecified functions map streams onto sets of streams, and they are isomorphic to relations on streams.
[…]
$f_{1}$ $\in T_{1}(f_{1},f_{2})$
$f_{2}$ $\in T_{2}(f_{1},f_{2})$
with two stream inclusions, which can easily be recast in the vector-valued form (1).
@@
If the stream transformer $T$ is deterministic then the stream inclusion (1) reduces to a system of stream equalities $f=T(f)$. Interacting systems of deterministic stream processors, for instance, are traditionally modeled as the least fixpoint of this equality, where $T$ is a Scott-continuous stream transformer in a complete partial order (Kahn, 1974). But such a denotational semantics does not extend naïvely to unbounded nondeterminism (Back, 1980; Staples and Nguyen, 1985; Pratt, 1984; Broy, 1986). One option are extensions of complete partial orders into power domains (Plotkin, 1976; Smyth, 1978), but there is no partial order on sets of streams that can be used as the approximation ordering for defining sets of least fixpoints of multi-valued functions on streams (Broy, 2023).
@@
We take a different approach by investigating under which conditions fixpoints exist for causal stream transformers with unconstrained nondeterminism, which is naturally modeled with multivalued maps. Here we distinguish between weakly and strongly causal stream transformers, whereby the latter notion also implies a strict, bounded delay for which the outputs are determined.
The main results for solving stream inclusions of the form (1) are as follows.
@@
  • For strongly causal stream transformers $T$ with nonempty, compact codomains, solutions $f\in T(f)$ are contained in the (unique) fixpoint $F={\mathrm{sp}}_{T}(F)$ for the strongest post of $T$ (Theorem 7.7). This latter fixpoint is obtained as the limit of a Picard iteration with an explicit quantitative bound on the approximation of each iteration.
@@
  • $f\in T(f)$ has a solution if $T$ is strongly causal, not the constant map to the empty set, and all the codomains of $T$ are closed in the topology induced by the prefix metric on streams (Theorem 7.12). In addition, we identify a slightly stronger condition than strong causality to establish the uniqueness of solutions of the stream inclusion (1). A fixpoint induction principle is derived to reason about these solutions.
  • $f\in T(f)$ has a solution if $T$ is strongly causal and all codomains of $T$ are nonempty and compact (Theorem 7.16).
@@
  • If the stream transformer $T$ is weakly causal then either $f\in T(f)$ has a solution or there exists a ball of streams with positive radius on which the distance, as measured by the prefix metric on streams, between input stream and corresponding output sets with respect to $T$ is, in a sense to be made precise, invariant  (Theorem 7.18).
@@
These developments are based on the correspondence between the logic-based notion of causality and the quantitative concept of contraction (and nonexpansion). More precisely, we show that a vector- and multivalued stream transformer is weakly causal if and only if it is nonexpansive, and it is strongly causal if and only if it is contractive for the prefix metric. This approach is motivated by recent work of Broy (Broy, 2023) on stream-based system design calculus for strongly causal transformers, which is based on Banach’s fixpoint principle to solve stream equations on deterministic stream transformers. We generalize this observation by (1) modeling nondeterministic and mutually dependent system components as conjunctions of causal stream inclusions, (2) establishing the equivalence of vector- and multivalued causal stream transformers with contraction in a spherically complete ultrametric space of streams based on the prefix distance of streams, and (3) applying multivalued fixpoint principles in this ultrametric stream space to obtain solutions of causal stream inclusions.
@@
First, Section 2 explains some basic concepts of distance and topological spaces that are used in this paper. The prefix distance between two streams is then measured in Section 3 using the longest common prefix. This results in a spherically complete ultrametric space of streams. The underlying concept of the prefix distance derives from Cantor sets, and it generalizes to arbitrary stream products and also to sets of streams via the induced Hausdorff distance. Next, Section 4 discusses the familiar concepts of stream transformers along with some horizontal and vertical composition operators. The focus in Section 5 is on causal stream transformers, whose outputs are determined by a finite history of inputs. For instance, causality is preserved under composition and refinement of stream transformers.
@@
In Section 6 we develop a metric-based characterization of causal stream transformers. More specifically, we show that a nondeterministic stream transformer is causal if and only if it is contractive with respect to the given prefix ultrametric on streams. This correspondence extends to a notion of Lipschitz contraction based on the Hausdorff distance between sets of streams, since stream transformers with nonempty compact images are Lipschitz contractive if and only if they are causal.
@@
Since the induced Hausdorff metric is a complete metric space on the set of nonempty compact sets, we obtain unique fixpoints in Section 7 for causal transformers such as the weakest pre and the strongest post set transformers. Moreover, there is a linearly, strictly decreasing upper bound of the prefix distance between this fixpoint and its approximation by the underlying Picard iteration, which suggests an anytime approximation algorithm. Furthermore, we formulate, in the spirit of Park’s lemma, induction principles for fixpoints of set transformers, and we show that the set of fixpoints of a multivalued stream transformer $T$ is contained in the fixpoint of the strongest post transformer for $T$.
@@
Additional fixpoint results in Section 7 construct solutions for the stream inclusion (1) by showing that every strongly causal vector- and multivalued stream transformer has a fixpoint as long as it is not the constant map to the empty set and its codomain is restricted to closed sets only. We also identify a slightly stronger condition for which this fixpoint is unique. The underlying Picard iteration enables us to derive an induction principle for reasoning about these fixpoints. In Section 7 we also state some immediate consequences of the correspondence of causality and Lipschitz contractivity for fixpoints of weakly causal maps. Section 8 discusses further consequences of the causality-contraction connection, and Section 9 concludes with a discussion on the relevance of these developments for the principled design of systems.
2. Preliminaries
We summarize a hodgepodge of concepts and notation for topological and metric spaces as they will be used in the remainder of this paper, with the intention of making the developments of this paper as self-contained as possible. Readers who are familiar with these fundamental concepts can proceed to the next section.
@@ 2. Preliminaries
For a metric space $(M,d)$ the sets $\mathrm{B}({x},{r})\,\coloneqq\,\{{y}\,\mid\,{d(x,y)<r}\}$ and $\mathrm{B}[{x},{r}]\,\coloneqq\,\{{y}\,\mid\,{d(x,y)\leq r}\}$ are called the open and closed balls of center $x$ and radius $r$, respectively. The family of open balls forms a base of neighborhoods for a uniquely determined Hausdorff topology on $M$, which is the topology induced by $d$ (on $M$). Open, closed, bounded, (dis)connected, convex, totally bounded (precompact), and compact sets are defined with respect to the metric-induced topology. The set of nonempty, closed, and bounded subsets of $M$, in particular, is denoted by ${\mathrm{CB}({M})}$, and ${\mathrm{Comp}({M})}$ is the set of nonempty compact subsets of $M$.
@@
The distance $d(a,B)$ of an element $a\in M$ to a nonempty set $B\subseteq M$ is defined by $d(a,B)=\inf_{b\in B}{d(a,b)}$. Clearly, $d(a,b)=d(a,\{b\})$ for all $a,b\in M$. For a bounded metric space $(M,d)$ the Hausdorff distance $\mathcal{H}_{d}(A,B)$ between two nonempty sets $A,B\subseteq M$ measures the ”longest path” to get from $A$ to $B$, or vice versa, from $B$ to $A$.
(2) $\mathcal{H}_{d}(A,B)$ $\,\coloneqq\,\max{(\sup_{x\in A}{(d(x,B))},\sup_{y\in B}{(d(y,A))})}\mbox{\@.}$
@@
A sequence $(x_{k})_{k\in\mathbb{N}}$ in $M$ is Cauchy if and only if for all $\varepsilon>0$ there exists $N\in\mathbb{N}$ such that $d(x_{n},x_{m})<\varepsilon$ for all $n,m\geq N$. Such a sequence $x_{k}$ converges to $x^{*}\in M$ if and only if for each neighborhood $U$ of $x$ there exists $N\in\mathbb{N}$ such that $x_{n}\in U$ for all $n\geq N$. Now, the space $M$ is Cauchy complete if every Cauchy sequence converges to some $x^{*}\in M$. Equivalently, $M$ is Cauchy complete if and only if the intersection of nested sequences of closed balls whose radius approaches to 00 are nonempty.
@@
A map $T:M\to M$ (in a metric space $M$) is contractive if there exists a constant $l$ with $10<l<$ 1 such that $d(T(x),T(y))\leq l\cdot d(x,y)$ for all $x,y\in M$. Traditionally, Banach’s contraction principle establishes that in a Cauchy complete metric space $(M,d)$ there is a unique fixpoint for every contracting $T:M\to M$ (Kelley, 1955). Starting with the Picard iteration $x_{n+1}=T(x_{n})$ for $n\in\mathbb{N}$ with $x_{0}$ arbitrary, one concludes from the contraction property with Lipschitz constant $10\leq l<$ 1, that $d(x_{n+1},x_{n})\leq l\cdot d(x_{n},x_{n-1})$, and, therefore, $d(x_{n+1},x_{n})\leq\nicefrac{{l^{n}}}{{1-l}}\cdot d(x_{0},x_{1})$. From this, one concludes that the sequence $(x_{n})_{n\in\mathbb{N}}$ is Cauchy, and, for completeness, that its limit $x^{*}\in M$ is a fixpoint. This fixpoint is unique, since we get $x^{*}=y^{*}$ for any two fixpoints from $d(x^{*},y^{*})=d(T(x^{*}),T(y^{*}))\leq l\cdot d(x^{*},y^{*})$.
@@
A map $T:M\to M$ is said to be shrinking if $d(T(x),T(y))<d(x,y)$ for all $x,y\in M$. A shrinking map $T$ need not have a fixpoint in a complete metric space.
(3) $d(x,y)$ $\leq\mathrm{max}({d(x,y)},{d(y,z)})\mbox{\@.}$
As a consequence of (3) the isosceles triangle principle
(4) $d(x,z)$ $=\mathrm{max}({d(x,y)},{d(y,z)})$
holds whenever $d(x,y)\neq d(y,z)$. Further immediate consequences of the strong triangle inequality are: (1) every point inside a ball is its center, that is, if $d(x,y)<r$ then $\mathrm{B}({x},{r})=\mathrm{B}({y},{r})$, (2) all balls of strictly positive radius $r$ are clopen, that is both open and closed, (3) if two balls are not disjoint then one is included in the other, (4) the distance $d(B_{1},B_{2})\,\coloneqq\,\inf_{x\in B_{1},y\in B_{2}}d(x,y)$ of two disjoint nonempty balls $B_{1}$,$B_{2}$ is obtained as the distance of two arbitrarily chosen elements $x\in B_{1}$, $y\in B_{2}$. (4) ultrametric spaces are totally disconnected, that is, every superset of a singleton set is disconnected, and (5) a sequence $(x_{k})_{k\in\mathbb{N}}$ in an ultrametric space is Cauchy if and only if $\lim_{k\to\infty}d(x_{k+1},x_{k})=0$. We will also make use of a generalized strong triangle inequality for ultrametric distances.
(5) $d(a,C)$ $\leq\max{(d(a,b),d(b,C))}\mbox{\@,}$
for all $a,b\in M$ and $\emptyset\neq C\subseteq M$.
Proposition 2.1.
An ultrametric space $(M,d)$ is spherically complete if $\,\cap\,_{{B\in\mathcal{C}}}B\neq\emptyset$ for every chain ${\mathcal{C}}$ of balls $B_{0}\supseteq B_{1}\supseteq B_{2}\supseteq\ldots$. From this definition it is clear that spherical completeness implies Cauchy completeness.
Example 2.2.
Proposition 2.3 (((Dieudonné, 2011), p. 59).
If $(M,d)$ is an ultrametric space then the ultrametric space $({\mathrm{Comp}({M})},\mathcal{H}_{d})$ is Cauchy complete.
For a spherically complete ultrametric space $M$, every strictly contracting $T:M\to M$ has a unique fixpoint in $M$ ((Petalas and Vidalis, 1993), Theorem 1). Moreover, if $T:M\to M$ is nonexpansive, that is $d(T(x),T(y))\leq d(x,y)$ for all $x,y\in M$, then either $T$ has at least one fixpoint or there exists a ball $B$ of radius $r>0$ such that $T:B\to B$ and for which $d(u,T(u))=r$ for each $u\in B$  ((Petalas and Vidalis, 1993), Theorem 2). Such a ball $B$ is said to be minimal $T$-invariant.
3. Streams
An $A$-valued stream in ${A}^{\omega}\,\coloneqq\,\mathbb{N}\to A$, for a given nonempty set $A$ of values, is an infinite sequence $(a_{k})_{k\in\mathbb{N}}$ with $a_{k}\in A$. Depending on the application context, streams are also referred to as discrete streams or signals, $\omega$-streams, $\omega$-sequences, or $\omega$-words. The generating function (Charalambides, 2018) of a stream $(a_{k})_{k\in\mathbb{N}}$ is a formal power series
(6) $\sum_{k\in\mathbb{N}}a_{k}X^{k}$
(7) $(\sum_{k\in\mathbb{N}}a_{k}X^{k})+(\sum_{k\in\mathbb{N}}b_{k}X^{k})$ $\,\coloneqq\,\sum_{k\in\mathbb{N}}(a_{k}+b_{k})X^{k}$
(8) $(\sum_{k\in\mathbb{N}}a_{k}X^{k})\cdot(\sum_{k\in\mathbb{N}}b_{k}X^{k})$ $\,\coloneqq\,\sum_{k\in\mathbb{N}}(\sum_{i=0}^{k}a_{i}b_{k-i})X^{k}$
If $A$ is a field, then ${A}^{\omega}\simeq{A}\llbracket{X}\rrbracket$ is a principal ideal domain with the ideal $(X)=X\cdot{A}^{\omega}$ the only nonzero maximal ideal. Moreover, for $A$ a field or a division ring, $({A}^{\omega},+,(a\cdot)_{a\in K})$ is a linear space, whereby the dot product $(a\cdot f)$, for $f\in{A}^{\omega}$, is defined by $\bar{a}\cdot f$. Unless stated otherwise, we assume the values $A$ in ${A}^{\omega}$ to be a field.
The valuation $v({f})$ of a stream $f\in{A}^{\omega}$ is the minimal $k\in\mathbb{N}$ such that $[X^{k}]f\neq 0$, if any exists; otherwise $v({f})\,\coloneqq\,\infty$. By construction,
(9) $|{f}|\,\coloneqq\,2^{-v({f})}\mbox{\@,}$